# Cross-Encoder Rediscovers a Semantic Variant of BM25

### v1

@@ Abstract.
Neural Ranking Models (NRMs) have rapidly advanced state-of-the-art performance on information retrieval tasks. In this work, we investigate a Cross-Encoder variant of MiniLM to determine which relevance features it computes and where they are stored. We find that it employs a semantic variant of the traditional BM25 in an interpretable manner, featuring localized components: (1) Transformer attention heads that compute soft term frequency while controlling for term saturation and document length effects, and (2) a low-rank component of its embedding matrix that encodes inverse document frequency information for the vocabulary. This suggests that the Cross-Encoder uses the same fundamental mechanisms as BM25, but further leverages their capacity to capture semantics for improved retrieval performance. The granular understanding lays the groundwork for model editing to enhance model transparency, addressing safety concerns, and improving scalability in training and real-world applications.
explainability, neural retrieval models, IR axioms, mechanistic interpretability
copyright: acmlicenseddoi: XXXXXXX.XXXXXXXconference: ; July 13–18, 2025; Padua, Italyccs: Information systems Retrieval models and rankingccs: Computing methodologies Causal reasoning and diagnostics
1. Introduction
@@
Information retrieval (IR) has long relied on foundational models like BM25, grounded in intuitive heuristics for relevance estimation. BM25 leverages term frequency (TF) and inverse document frequency (IDF) to rank documents effectively, achieving strong performance across various tasks. Its simplicity and interpretability have made it a cornerstone of traditional IR systems. Inspired by BM25’s success, earlier neural IR models (e.g., MatchPyramid (Pang et al., 2016), DRMM (Guo et al., 2016)) were purposefully designed to emulate BM25’s principles. These models incorporated explicit components for semantic TF and IDF computations, blending neural architectures with established IR heuristics to improve relevance estimation.
@@
However, the advent of transformer-based models has revolutionized the field of IR. These models, trained end-to-end on raw data, excel at extracting context-dependent semantic signals for ranking tasks. By leveraging the multi-attention mechanism and vast parameter spaces, transformers (Vaswani et al., 2023) capture nuanced relationships between query and document terms that go beyond traditional heuristic-based approaches. Despite their superior performance, these models come with significant trade-offs: their complexity and lack of interpretability make it challenging to understand how they assess relevance. This raises fundamental questions: how do these models assess relevance? Do they rely on established IR principles such as TF and IDF, or do they operate as vast, unstructured statistical approximators?
@@
In this work, we build upon previous correlational studies by employing mechanistic interpretability (Nanda, 2022; Elhage et al., 2021; Olsson et al., 2022; Meng et al., 2024; Wang et al., 2022; Pearl, 2022) methods to address these questions. These techniques have further advanced our understanding of the Transformer-based architectures (Vaswani et al., 2023) that underpin many modern NLP systems (Olsson et al., 2022; Geiger et al., 2021; Meng et al., 2024; Wang et al., 2022; Vig et al., 2020) and in IR, have been used to identify components responsible for encoding term frequency signals in bi-encoders (Chen et al., 2024).
In this paper, we deeply analyze a cross-encoder and validate prior research suggesting the presence of a BM25-like signal by causally demonstrating how a semantic variant of BM25 is precisely implemented within the components of a transformer architecture.
Specifically, our main findings are:
  • 4.14.3) We use path patching to identify attention heads that compute and process BM25-like components, including a semantic version of TF (soft-TF), term saturation, and document length. These heads send this information to another group of heads (Relevance Scoring Heads), which we hypothesize compute a BM25-style function by aggregating all these relevance signals.
  • 4.4): We find evidence that IDF information exists in a dominant low-rank vector of the embedding matrix and can be manipulated to control term importance.
  • 5): We confirm that Relevance Scoring Heads implement a BM25-style computation by defining a linear approximation of the cross-encoder’s relevance computation and evaluating its ability to reconstruct scores, confirming that our circuit understanding captures the core mechanism of relevance computation.
Table 1. Decomposition of BM25 into components and associated axioms. Note: Since IDF is a distinct component in BM25 (unlike term saturation and document length which are tied to TF), we use alternative methods to isolate this component (§4.4).
@@
TF (soft-TF) TFC1 (Fang et al., 2004): Prefer documents with more query term occurrences. Given a baseline document, we perturb it by appending one more selected query term. b: Quebec is a small city in Canada.
@@
STMC1 (Fang and Zhai, 2006): Prefer documents with semantically similar terms. Given a baseline document, we append a semantically similar term (the term with highest cross-encoder embedding similarity among 20 GPT-4o-generated candidates.) b: Quebec is a small city in Canada.
@@
IDF TDC (Fang et al., 2004): Prefer documents with more discriminative query terms. N/A N/A
@@
k TFC2 (Fang et al., 2004): Additional occurrences yield smaller improvements. We create a baseline document using GPT-4o to create five relevant sentences starting with selected query term’s pronoun and perturb by incrementally restoring the term. b: It is in Canada. It is a fun city.
p1: Quebec is in Canada. It is a fun city.
p2: Quebec is in Canada. Quebec is a fun city.
@@
b LNC1 (Fang et al., 2004): Penalize longer documents for non-relevant terms. Given a baseline document, we create five perturbations by sequentially appending an increasing number of random sentences from a random query in the base dataset, adding non-relevant information. b: Quebec is in Canada.
p1: Quebec is in Canada. I love papaya ice-cream at Abby’s.
p2: Quebec is in Canada. I love papaya ice-cream at Abby’s. Forever and Ever is a song by Julia.
By leveraging path patching for a sequential tracing of subcomponents on the cross-encoder’s term-matching path, we reverse-engineer a dominant portion of cross-encoder’s pathway to compute relevance. Such granular understanding reveals potential for model-editing, potentially enabling model performance enhance, controllability and personalization, and mitigate biases for NRMs.
2. Related Work
2.1. Interpreting Neural IR Models
In contrast to early neural IR models (e.g., MatchPyramid (Pang et al., 2016), DRMM (Guo et al., 2016)) that were explicitly designed to be neural term matching models, recent pre-trained transformer-based ranking models achieve significant performance gains by capturing nuanced query-document interactions and forming rich contextual representations (Trabelsi et al., 2021; Guo et al., 2019). Although not explicitly designed to encode relevance concepts, these models have been found to learn relevance concepts such as TF (Chen et al., 2024), IDF (Choi et al., 2022), and even combinations of them such as BM25 (Wang et al., 2021; Rau and Kamps, 2022; Yates et al., 2021; MacAvaney et al., 2022). To better understand how these relevance signals are represented and utilized, researchers have turned to IR axioms (Câmara and Hauff, 2020; Rennings et al., 2019a; Völske et al., 2021) and constructed axiom-specific diagnostic datasets to test how well models rank document pairs in accordance with axiomatic constraints (Rennings et al., 2019b). Many studies also rely on correlational methods such as probing (Choi et al., 2022; Zhan et al., 2020; Formal et al., 2021, 2022; MacAvaney et al., 2022) to diagnose model behavior and assess their adherence to desired relevance properties. However, the underlying mechanisms of how these concepts are encoded remain unclear. This work aims to identify the specific components involved in this process and explain how they interact to determine relevance.
2.2. Mechanistic Interpretability
To better understand how information is encoded in models, mechanistic interpretability (Nanda, 2022; Elhage et al., 2021; Olsson et al., 2022; Meng et al., 2024; Wang et al., 2022; Pearl, 2022) aims to reverse-engineer model behavior by analyzing the causal effects of components, providing stronger causal evidence than probing (Vig et al., 2020; Meng et al., 2024). Gaining traction in NLP, these methods have been used to uncover how LLMs perform tasks, such as indirect object identification (Meng et al., 2024) and greater-than computation (Hanna et al., 2024), and have recently been applied to IR to identify attention heads encoding term frequency signals (Chen et al., 2024).
Mechanistic interpretability methods rely on causal interventions, such as activation patching (or causal mediation analysis) where activations are swapped between input pairs to observe their effects on model behavior and localize specific mechanisms. Activation patching provides a solid foundation for initial analyses, allowing researchers to identify and isolate specific mechanisms within a model (Zhang and Nanda, 2023; Heimersheim and Nanda, 2024). For example, Chen et al. (2024) use activation patching with IR axioms to identify attention heads that encode a term frequency signal in a bi-encoder. However, activation patching has limitations as it only evaluates the indirect effect components have on the output. In contrast, path patching allows us to trace the computational path (circuit) from input to output, providing insights into how components sequentially interact (Wang et al., 2022; Goldowsky-Dill et al., 2023). In this work, we leverage path patching to reverse-engineer the computational path through which a NRM uses to compute relevance.
3. Methodology
Figure 1. Path patching methodology. Left: There are four forward passes: (1,2) Run model on baseline and perturbed inputs and cache activations. (3) To measure the effect of an upstream sender component (S) on a downstream receiver component (R), run the model on the baseline input, patch in S, freeze all other components, and cache the activation of R. (4) Run the model on the baseline input and patch in R. Right: Alternative visualization of step (3) using the residual stream. By allowing only the downstream receiver R to be recomputed when the sender S is patched, we effectively isolate the direct path from S to R, while preserving all other paths to R as they were in the baseline run.
In this section, we explain how we construct diagnostic datasets and use path patching to track how the model implements a semantic version of BM25.
3.1. Diagnostic Dataset Construction
Mapping BM25 components to IR axioms. IR axioms are derived from the same retrieval heuristics BM25 is composed of: TF, IDF, term saturation, and document length (Fang et al., 2004). Table 1 outlines the mapping from BM25 components to axioms and provides their formal definitions. Note that for TF, we extend this analysis of term matching to include semantic matches, where semantically similar terms contribute to the relevance score, based on the intuition about the semantic understanding capabilities of Transformer-based models (Trabelsi et al., 2021). We refer to this extension as soft-TF and utilize both TFC1 and STMC1 as the corresponding axioms.
@@ 3.1. Diagnostic Dataset Construction
Constructing minimal-pair-based diagnostic datasets. We construct a diagnostic dataset for each axiomatic component by perturbing documents to create input pairs. Our starting point is the base dataset from Chen et al. (Chen et al., 2024), who create a diagnostic dataset for analyzing TFC1 using MS-MARCO (Nguyen et al., 2016), consisting of 10k query-document pairs. To create the diagnostic dataset for each axiom, we perturb each query-document pair in the base dataset in accordance with the formal axiomatic definition (Table 1). Each perturbed sample differs from the original only slightly, with the perturbation introducing an additional signal corresponding to the relevance concept we aim to analyze. Our perturbation strategies and examples of input pairs are shown in Table 1.
3.2. Path Patching
The intuition of path patching relies on the Transformer (Vaswani et al., 2023) architecture, where stacks of multi-headed attention and multi-layer perceptron (MLP) layers are connected by residual connections. These residual connections are additive functions, which essentially allow components to “read” and “write” information to a common residual stream.  (Team, 2021). The goal of path patching is to isolate which components “send” and “receive” information through the residual stream, and what information is being transmitted.
Path patching (Wang et al., 2022; Goldowsky-Dill et al., 2023) involves using minimal input pairs (we denote as baseline b and perturbed p), that differ only by a target behavioral signal to be evaluated (e.g., axiomatic property). Activations (e.g., residual stream output, attention head or layer output, MLP output) are iteratively copied from the perturbed input into the corresponding location in the baseline input to isolate the impact of that component on downstream computations. In this work, we focus on analyzing the roles of attention heads in relevance computation and leave analysis of MLPs for future investigation.
The effect of a patch is measured by the difference in logits (which in the case of cross-encoders, is equivalent to the difference in relevance scores). This algorithm is then iteratively repeated for each important upstream component. For more information, we refer the reader to Wang et al. (Wang et al., 2022) and Goldowsky-Dill et al. (Goldowsky-Dill et al., 2023).
3.3. Model
We analyze ms-marco-MiniLM-L-12-v2 (Face, 2025) for our BERT-based cross-encoder, a model trained on the MS MARCO dataset, and chosen for its high evaluation performance the MS MARCO Passage Re-Ranking  (Nguyen et al., 2016) and TREC Deep Learning 2019 datasets (Craswell et al., 2020). The cross-encoder processes input x, consisting a query $q$ and a document $d$ in the format: <CLS> q1, q2, ... <SEP> d1, d2, ... <SEP>. After all transformer layers, the <CLS> token vector is passed to a classifier to produce the logits.
4. Semantic Scoring Circuit
Previous work in mechanistic interpretability has identified groups of attention heads whose behaviors generalize across tasks. For example, ”induction heads,” ”previous token heads,” and ”duplicate token heads” (Olsson et al., 2022) have been shown to play a role in various NLP tasks by completing sub-tasks within larger task-specific circuits. Identifying these heads is therefore crucial. Following the convention of naming head groups based on their mechanisms, we assign names to key components according to their specialized roles in document relevance computation. Figure 2 illustrates the uncovered relevance circuit, with details provided below:
  • Matching Heads generate a Matching Score by summing soft-TF values for each query term, while accounting for term saturation and document length effects.
  • Contextual Query Representation Heads distribute soft-TF information of the higher-IDF query tokens among all query tokens to strengthen the higher-IDF tokens’ representations.
@@ 4. Semantic Scoring Circuit
  • IDF of vocabulary terms is stored in a dominant low-rank vector of the model’s Embedding Matrix.
  • Previously listed components send the four BM25 signals to Relevance Scoring Heads, which we hypothesize are summing Matching Score weighted by IDF in a BM25 manner.
Figure 2. We find that the cross-encoder rediscovers a semantic variant of BM25. Specifically, Matching Heads, Contextual Query Representation Heads, and the embedding matrix compute, process and send BM25-like components’ information to the Relevance Scoring Heads which finally compute the relevance score in a BM25-like manner.
As described in §3, we analyze the critical soft-TF component using path patching, tracing heads from ”back” to ”end” to align with its iterative nature. We first identify attention heads directly influencing logits and hypothesize they perform BM25-like computations by summing soft-TF weighted by IDF. Tracing further, we uncover Contextual Query Representation Heads and Matching Heads where soft-TF is originally computed. Using other diagnostic datasets, we find term saturation and document length signals intertwine with soft-TF at the Matching Heads. After tracking soft-TF, we apply SVD and intervention and localize IDF values in the first low-rank vector of the Embedding. We localize all BM25-like components in this section and verify that Relevance Scoring Heads perform BM25-like computations in the next section.
4.1. Relevance Scoring Heads
Figure 3. Path patching from individual attention heads to final logits reveals 10.1, 10.4, 10.7, 10.10 as the most critcal head in sending soft-TF information to [CLS] for ranking score computation for both TFC1 and STMC1 diagnostic datasets. Notably, these critical heads are shared between TFC1 and STMC1, exhibiting very similar importance values.
Information Flow. To identify components directly transmitting soft-TF signals to the relevance scores, we path patch to the logits using the TFC1 and STMC1 diagnostic datasets.
Figure 3 highlights the most significant heads (i.e., those causing a $>30\%$ increase in ranking score) for the soft-TF-related axioms: 10.1 (Layer 10, Head 1), 10.4, 10.7, and 10.10. The high correlation in patching importance values (corr = 0.9996, p ¡ 0.001) and the fact that the four critical heads are shared between TFC1 and STMC1 suggests that the model treats exact and semantic term matches similarly.

### v2

@@ Abstract
Neural Ranking Models (NRMs) have rapidly advanced state-of-the-art performance on information retrieval tasks. In this work, we investigate a Cross-Encoder variant of MiniLM to determine which relevance features it computes and where they are stored. We find that it employs a semantic variant of the traditional BM25 in an interpretable manner, featuring localized components: (1) Transformer attention heads that compute soft term frequency while controlling for term saturation and document length effects, and (2) a low-rank component of its embedding matrix that encodes inverse document frequency information for the vocabulary. This suggests that the Cross-Encoder uses the same fundamental mechanisms as BM25, but further leverages their capacity to capture semantics for improved retrieval performance. The granular understanding lays the groundwork for model editing to enhance model transparency, addressing safety concerns, and improving scalability in training and real-world applications. [^1] All code and resources are available at: https://github.com/mlu108/CrossEncoderBM25
Cross-Encoder Rediscovers a Semantic Variant of BM25
Meng Lu Brown University meng_lu@brown.edu                        Catherine Chen Brown University catherine_s_chen@brown.edu
Carsten Eickhoff University of Tübingen carsten.eickhoff@uni-tuebingen.de
1 Introduction
@@
Information retrieval (IR) is the subfield of NLP that aims to rank a collection of documents by their relevance to a specific query. Traditional ranking strategies have long relied on probabilistic models grounded in intuitive heuristics, like BM25 (Robertson and Zaragoza, 2009), to estimate relevance. BM25 leverages term frequency (TF) and inverse document frequency (IDF) to rank documents effectively, achieving strong performance across various tasks. Its simplicity and inherent interpretability have made it a cornerstone of traditional IR systems. Inspired by BM25’s success, earlier neural IR models (Pang et al., 2016; Guo et al., 2016) were purposefully designed to emulate BM25’s principles. These models incorporate explicit components for semantic TF and IDF computations, blending neural architectures with established IR heuristics to improve relevance estimation.
@@
However, the advent of transformer-based models revolutionized the field of IR. These models, trained end-to-end on large numbers of query-document pairs (Nguyen et al., 2016; Thakur et al., 2021), excel at extracting context-dependent semantic signals for ranking tasks. By leveraging multi-headed attention and vast parameter spaces, transformers Vaswani et al. (2017) capture nuanced relationships between query and document terms that go beyond traditional heuristic-based approaches. Despite their superior performance, these models come with significant trade-offs: their complexity and lack of interpretability make it challenging to understand their internal mechanisms. This raises fundamental questions: how do these models assess relevance? Do they rely on established IR principles such as TF and IDF, or do they draw on entirely different expressions of relevance?
@@
In this work, we build upon previous correlational studies (Choi et al., 2022; Zhan et al., 2020; Formal et al., 2021, 2022; MacAvaney et al., 2022) by employing mechanistic interpretability methods (Nanda, 2022; Elhage et al., 2021; Olsson et al., 2022; Meng et al., 2024; Wang et al., 2022; Pearl, 2022) to address these questions.
Concretely, we analyze a BERT-based IR model (i.e., a cross-encoder) to isolate multiple relevance signals beyond exact match TF. Our main findings are: (1) We identify attention heads that extract and process BM25-like components, including a semantic version of term frequency (soft-TF), term saturation, and document length (§4.14.3). (2) We find evidence that IDF information largely already exists in the largest singular value of the embedding matrix and can be manipulated to later control term importance (§4.4). (3) We confirm that heads in later layers aggregate all these relevance signals in a BM25-style manner by defining a linear approximation of the hypothesized relevance computation and evaluating its ability to reconstruct cross-encoder’s predicted scores, confirming that our circuit captures the core mechanism of relevance computation (§5).
Overall, our mechanistic analysis uncovers insights on how IR models determine relevance, paving the way for targeted model editing to boost performance, enable personalization, and mitigate bias in transformer-based IR.
2 Background and Related Work
2.1 Axioms and BM25
@@
TF (soft-TF) TFC1 Fang et al. (2004): Prefer documents with more query term occurrences. Given a baseline document, we perturb it by appending one more selected query term. b: Quebec is a small city in Canada.
@@
STMC1 Fang and Zhai (2006): Prefer documents with semantically similar terms. Given a baseline document, we append a semantically similar term - the synonym with highest embedding cosine similarity with query term from 20 candidates generated by GPT-4o. b: Quebec is a small city in Canada.
@@
IDF TDC Fang et al. (2004): Prefer documents with more discriminative query terms. N/A N/A
@@
k TFC2 Fang et al. (2004): Additional occurrences yield smaller improvements. We create a baseline document using GPT-4o to create five relevant sentences starting with selected query term’s pronoun and perturb by incrementally restoring the term. b: It is in Canada. It is fun.
p1: Quebec is in Canada. It is fun.
p2: Quebec is in Canada. Quebec is fun.
@@
b LNC1 Fang et al. (2004): Penalize longer documents for non-relevant terms. Given a baseline document, we create five perturbations by sequentially appending an increasing number of random sentences from a non-relevant query in the base dataset. b: Quebec is in Canada.
p1: Quebec is in Canada. Road not taken is written by Frost.
p2: Quebec is in Canada. Road not taken is written by Frost. Happiness is a Butterfly.
Table 1: Mapping of BM25 components to IR axioms. Note: Since IDF is a distinct component in BM25 (unlike term saturation and document length, which are tied to TF), we use alternative methods to isolate it (§4.4).
Axiomatic IR constructs formal desiderata, or axioms, outlining specific properties that an effective ranking model should satisfy (Bruza and Huibers, 1994). For example, the TFC1 axiom (Fang et al., 2004) states that documents that contain more query terms should be ranked higher, and the TDC axiom (Fang et al., 2004) states that documents containing query terms with higher inverse document frequency (IDF) should receive higher scores. These axioms provide a theoretical foundation for understanding and developing ranking functions by formalizing human-interpretable and intuitive notions of relevance into mathematical constraints.
BM25 is a widely used probabilistic ranking function that exemplifies these axiomatic principles by ranking documents based on term frequency (TF), inverse document frequency (IDF), term saturation, and document length. It is defined as:
$\sum_{t\in q}\text{IDF}(t)\cdot\frac{\text{TF}(t,d)\cdot(k_{1}+1)}{\text{TF}(t,d)+k_{1}\cdot(1-b+b\cdot\frac{|d|}{\text{avgdl}})}$ (1)
where $q$ is the query, $d$ is the document, $t$ is a query term, $k_{1}$ and $b$ are hyperparameters controlling term saturation and length normalization, respectively. $|d|$ is the length of document $d$, and avgdl is the average document length in the corpus.
Intuitively, BM25 computes a relevance score by multiplying each query term’s term frequency (TF) in the document with its inverse document frequency (IDF), capturing both how relevant and how informative the term is. It then sums these scores across all query terms with two adjustments: diminishing returns for repeated terms (term saturation via $k_{1}$), and downweighting for longer documents (length normalization via $|d|$), which prevents overly long documents with additional query term occurrences from being unfairly favored. As shown in Table 1, each component of BM25 aligns with a specific axiom. For instance, TFC1 reflects the TF factor, which favors documents containing more query terms. Framing BM25 in terms of axioms allows us to mechanistically analyze whether and how neural IR models internalize these principles when estimating relevance.
2.2 Mechanistic Interpretability for IR
Despite not being explicitly trained to encode traditional relevance concepts, previous work provides correlational evidence suggesting that BERT-based IR models encode BM25-like information  (Wang et al., 2021; Rau and Kamps, 2022; Yates et al., 2021; MacAvaney et al., 2022; Choi et al., 2022; Zhan et al., 2020; Formal et al., 2021, 2022).
To move beyond correlation and gain a causal understanding of how relevance is computed, we turn to mechanistic interpretability (Nanda, 2022; Elhage et al., 2021; Pearl, 2022), which has been instrumental in uncovering how transformer-based NLP models perform certain tasks, such as indirect object identification Meng et al. (2024) and greater-than computation Hanna et al. (2024).
In the context of IR, Chen et al. (2024) apply activation patching (Vig et al., 2020; Meng et al., 2024) to identify attention heads responsible for exact term matching. We build on this work by decomposing BM25 into intuitive IR axioms and constructing diagnostic datasets to localize each component. Rather than focusing solely on exact matches, we extend the analysis to semantic matches and trace how these components interact to compute the final relevance score using path patching (Wang et al., 2022; Goldowsky-Dill et al., 2023). This work is the first to causally uncover an internal circuit for the relevance estimation mechanism of neural IR architectures.
3 Methodology
3.1 Model
While many early neural models were purposefully designed to emulate BM25’s principles, transformer-based models have revolutionized the field of IR. Among them, cross-encoders represent a specific architectural approach to neural ranking models that process query-document pairs jointly.
Given input x in the format: <CLS> query <SEP> document <SEP>, the model is trained to optimize a binary classification task where the <CLS> token is passed to a classifier to determine whether the provided query is relevant to the document. In this work, we choose to examine ms-marco-MiniLM-L-12-v2 (Face, 2025), a BERT-based cross-encoder, for its high performance on common IR benchmarks (Nguyen et al., 2016; Craswell et al., 2020).
3.2 Diagnostic Datasets Construction
@@
To investigate if and how the model implements BM25, we map BM25 components to IR axioms (§2.1) and construct diagnostic datasets that isolate individual axiomatic components. We build off of Chen et al. (2024), who create a diagnostic dataset for analyzing TFC1 using MS-MARCO (Nguyen et al., 2016), consisting of 10k web search query-document pairs. To create a diagnostic dataset for the remaining axioms, we perturb each query-document pair in this base dataset following their formal axiomatic definition (Table 1). Each perturbed sample differs minimally from the original, with the perturbation introducing an additional signal corresponding to the axiom. Our perturbation strategies and examples of input pairs are shown in Table 1. Additionally, we expand the analysis on TF to soft-TF to include semantic matches, motivated by findings that neural models often score documents highly even in the absence of exact lexical matches (Rau and Kamps, 2022).
3.3 Path Patching on Diagnostic Datasets
We first begin by tracking term matching, the core feature of classical IR models like BM25, by applying path patching on minimal pairs in TFC1 and STMC1 diagnostic datasets to uncover the soft-TF mechanism. For other components in BM25, we use these datasets to further inspect how they interact with the main soft-TF circuit.
Specifically, we isolate soft-TF by selecting a query, a baseline document b, and a perturbed document p, where p introduces an additional semantic or lexical term frequency signal. We then substitute activations from p into b at specific model components, such as individual attention heads or MLP layers. If this substitution causes the model’s relevance score for b to shift toward that of p, we interpret it as causal evidence that the patched component encodes a term matching signal.
The path patching algorithm proceeds iteratively in a backward fashion: it first identifies components directly responsible for changes in the output logits, then recursively traces upstream to uncover the full set of causally relevant components. This backward search reveals the components through which specific information, introduced by the perturbation, flows (more details in Appendix §A).
4 Semantic Scoring Circuit
Figure 1: Overview of relevance mechanisms in the model.
In the course of this section, we will identify and localize the following components in the model’s Semantic Scoring Circuit (Figure 1):
  • Matching Heads locate exact and semantic term matches in the document for each query term, while also encoding term saturation and document length signals.
  • Query Contextualization Heads distribute the matching signal from higher-IDF query tokens to all query tokens.
@@ 4 Semantic Scoring Circuit
  • IDF of each term is stored in a dominant low-rank vector of the model’s embedding matrix.
  • Relevance Scoring Heads aggregate query-term importance by combining the matching signal (soft-TF) and IDF for each query term in a manner similar to BM25.
As described in §3, we begin by uncovering the core soft-TF mechanism via a “backward pass” of the model: path patching to the output logits, then progressively patching to earlier components to sequentially uncover important heads that compute and propagate soft-TF signals.
After localizing the soft-TF mechanism, we apply Singular Value Decomposition (SVD) on the embeddings to investigate IDF. After isolating all BM25-like components in this section, we verify that Relevance Scoring Heads perform a BM25-like computation in §5.
4.1 Relevance Scoring Heads
Figure 2: Path patching identifies heads 10.1, 10.4, 10.7, and 10.10 as the most important carriers of soft-TF signals to [CLS] on both TFC1 and STMC1, with similar patching effects.
Information Flow. To identify the components that directly transmit soft-TF signals to the relevance scores, we path patch to the logits using the TFC1 and STMC1 diagnostic datasets. We observe a high Pearson correlation between the patching effects of TFC1 and STMC1 (r = 0.99, p < 0.001), indicating that the same heads produce similar effects on delivering soft-TF information into the final logits (see Appendix 9 for detailed patching results). Figure 2 displays the heads with the highest patching effects, which are consistent across both TFC1 and STMC1: 10.1 (Layer 10, Head 1), 10.4, 10.7, and 10.10. Notably, these heads contribute a larger change in relevance scores for TFC1 compared to STMC1, suggesting that the model may prioritize exact matches over semantic ones.